# Erasure Minesweeper: exploring hybrid-erasure surface code architectures for efficient

### v1

@@ Abstract
Dual-rail erasure qubits can substantially improve the efficiency of quantum error correction, allowing lower error rates to be achieved with fewer qubits, but each erasure qubit requires 3×3$\times$3 × more transmons to implement compared to standard qubits. In this work, we introduce a hybrid-erasure architecture for surface code error correction where a carefully chosen subset of qubits is designated as erasure qubits while the rest remain standard. Through code-capacity analysis and circuit-level simulations, we show that a hybrid-erasure architecture can boost the performance of the surface code—much like how a game of Minesweeper becomes easier once a few squares are revealed—while using fewer resources than a full-erasure architecture. We study strategies for the allocation and placement of erasure qubits through analysis and simulations. We then use the hybrid-erasure architecture to explore the trade-offs between per-qubit cost and key logical performance metrics such as threshold and effective distance in surface code error correction. Our results show that the strategic introduction of dual-rail erasure qubits in a transmon architecture can enhance the logical performance of surface codes for a fixed transmon budget, particularly for near-term-relevant transmon counts and logical error rates.
@@ I Introduction
Figure 1: Viewing a hybrid-erasure architecture as a game of Minesweeper. Top: In a standard surface code, we receive syndrome bits (red and blue 1s) indicating that an error string terminates nearby. The decoder must find a valid explanation (red and blue flags) based only on this information. If the decoder guesses wrong, a logical error can occur. Bottom: When some data qubits are erasure qubits, we gain extra information before decoding (black bombs and green checks on left) that help us decide which error string to predict, avoiding logical errors.
@@
While typical noise models used to simulate and evaluate QEC assume that physical one- and two-qubit errors are random, uniformly distributed Pauli operators, significant gains in QEC efficiency can be found under alternative noise models [8, 9, 10, 11, 12, 13]. For example, under a biased noise model where the relative strengths of Pauli X and Z errors are orders of magnitude different, the XZZX surface code can drastically improve the noise threshold compared to the standard surface code [8]. Another type of non-standard error model is heralded erasure, in which the information in a physical qubit is completely lost, but the time and location of the error are known. A surface code made up of erasure qubits, in which the primary error mechanism is heralded erasure, can have a higher threshold and, crucially, twice the effective distance: a surface code patch with $d\times d$ data qubits can correct $d-1$ heralded erasure errors on ideal erasure qubits, but only $\lfloor(d-1)/2\rfloor$ errors under a standard Pauli error model. This gives erasure qubits a significant scaling advantage over standard unbiased qubits, leading to significant interest in the proposal [14, 15, 16, 17] and realization [18, 19, 20, 21, 22, 23, 24, 25, 26] of erasure qubits in various hardware modalities.
Dual-rail erasure qubits in superconducting transmons [16, 15] are a particularly promising approach to implementing erasure qubits, with several recent experimental realizations [22, 20, 26]. A dual-rail qubit is encoded in two transmons such that a relaxation error (the dominant error in individual transmons) in either transmon takes the system out of the qubit space; an erasure check can be performed regularly to determine whether this has happened, and if so, the qubit is reset and the information is sent to the decoder. Some dual-rail implementations require a third transmon to perform this erasure check, bringing the relative cost of a single dual-rail erasure qubit to 3×3$\times$3 × that of a standard qubit.
@@ II-B Erasure qubits
[…]
@@ III Analysis of a hybrid-erasure surface code
  • $P\subset[d^{2}]$ specifies the subset of data qubits (indexed by $[d^{2}]=\{1,..,d^{2}\}$) designated as erasures.
@@ III-A Summary of results
Figure 4: Examples of hybrid-erasure architecture. (a) $\mathcal{A}(5,0.6,P_{r})$ where the placement $P_{r}=\{1,2,5,...,22\}$ is sampled at random. (b) $\mathcal{A}(7,0.57,P^{*})$. Since $f_{e}\times d^{2}=27.93$, only $27$ erasures are allocated, with $24$ erasures placed in $2$ rows/columns, and the remaining $3$ placed greedily, all as close to the center as possible.
@@ III-B The repetition code: a toy model
Figure 5: Error correction on the repetition code with heralded erasure noise. (a) We consider a distance-$7$ repetition code with $k=3$ erasure qubits, at indices $q_{2},q_{3},q_{6}$. (b) Suppose we encounter $3$ physical $X$ errors, triggering a detection event on $q_{3}$. (c) A physical error on $q_{2}$ may trigger the detection event, but does not flip any syndromes; this could be due to the deplorizing error being projected onto the code space as a $Z$ error or an identity, each with probability $25\%$. (d) There are two ways to match the syndromes. Since there’s no detection even on $q_{6}$, we know the error must be the error chain not containing $q_{6}$.
@@
×$\times p^{l_{d}}(1-p)^{d-k-l_{d}}(\frac{1}{2})^{k}$
×$\times\chi[2l_{d}<d-k]$
@@ III-C Impact of erasure placement on logical error rate
We show in Figure 6 a histogram of how frequently each qubit in a distance $7$ surface code is contained in a logical error. Immediately, we observe that the qubits in the middle are included in up to 3×3$\times$3 × more paths than on the edge. We derive our closest-to-center placement strategy from this observation. Secondly, a more subtle point is that, since each lattice path either must traverse all columns or all rows, filling an additional row and column with erasures necessarily guarantees that each lattice path encounters an additional erasure qubit. The effective distance of the code will be guaranteed to increase by the same analysis given in Section III-B. It is important to note that, while in general we should place erasures where they encounter more traversing paths (in the center), we should not prioritize this over filling additional columns/rows, since, to leading order, the logical error performance is constricted by the error with the shortest fault distance.
@@ IV-A Noise model
Initialization error $p$ $2p$
Readout error $p$ $2p$
@@
Figure 8: Effective distance $d_{\text{eff}}$ of surface codes at various erasure qubit fractions. Optimized placement yields higher effective distance for intermediate erasure fractions. Lines show linear interpolation between $(d+1)/2$ and $d$.
@@ IV-B Simulation and decoding
We perform circuit-level simulations using Stim[47]. Errors are added upon initialization, readout, single-qubit gates, two-qubit gates, and erasure checks according to the noise model in Table I. One- and two-qubit gate errors on erasure qubits are modeled as heralded detections followed by a replacement of the qubit(s) with a maximally mixed state. The resulting circuit-level effect is an erasure flag along with $\{I,X,Y,Z\}$ errors with probability $\frac{1}{4}$ for each affected qubit.
@@
where $p_{e}$ is the probability of error $e$ derived from the noise model in Table I. See Figure 3 for an example.
@@ IV-C Calculating effective distance and threshold
Figure 9: Extracted surface code thresholds for random and optimized placement for different erasure fractions. Optimized placement leads to more consistent improvements in threshold.
@@ V Evaluating hybrid-erasure architectures
Figure 10: Achievable logical error rate for hybrid-erasure surface codes as a function of transmon count, for several fixed physical error rates. Hybrid architectures outperform all-standard or all-erasure architectures for certain transmon budgets.
Figure 11: Calculating expected fabrication yield for a chip containing the number of transmons needed for each surface code instance (see Figure 10). Assuming a defect rate of $10^{-3}$ per transmon.
@@ V-A Effective distance and threshold
Figure 12: Scaling of costs in Figure 10 for larger systems using an all erasure architecture, a no erasure architecture, or a hybrid erasure architecture. All hybrid points sweep over and select the best erasure fraction. Shaded region represents area of possible hybrid erasure values in between analytical upper bound and approximate observed circuit-level lower bound. (Top) Minimum possible logical error rate for fixed transmon cost. Background gradient indicates resulting single chip yield for a defect rate of $10^{-3}$. (Bottom) Minimum possible number of transmons for a target logical error rate.
@@
Figure 9 shows the extracted surface code thresholds for varying erasure qubit fraction, similarly showing a steady improvement as we add more erasure qubits. Again, a random placement of erasure qubits significantly underperforms compared to the optimized heuristic.
@@ V-B Comparing transmon counts
In this context, we examine the transmon cost of hybrid-erasure architectures in Figure 10 for several fixed physical error rates. These can be viewed as vertical slices through Figure 7, with transmon count calculated as a function of code distance and erasure fraction. For a fixed code distance and physical error rate, we see the expected increase in transmon count and improvement in logical error rate as erasure fraction increases. For certain fixed transmon budgets, e.g. from 70 to 100 or from 150 to 190 transmons, a hybrid-erasure design gives a better logical error rate than can be achieved with a full-erasure design.
@@ V-C Scaling to larger transmon counts
Although Figures 10 and 11 help analyze costs for near-term system sizes, we are also interested in the impact of hybrid erasure architectures for larger systems. To estimate the performance of the hybrid erasure architecture in higher distances and lower logical error rates, we use the empirical thresholds and a projected $d_{\text{eff}}$ bounded between two estimates. As an optimistic upper bound, we may expect the erasure fraction to behave as in the repetition code described in Section III-B. We observe that the fit becomes more accurate as we increase the distance of the surface code. The analytical lower-bound is derived using Equation 2. In Figure 12 , we compare a full-erasure architecture, a no-erasure architecture, and a hybrid architecture for surface code sizes up to 1,000 transmons and down to logical error rates of $10^{-15}$, where the shaded regions represent the parameter regimes where a hybrid architecture may potentially outperform the standard approaches. In the top plot of Figure 12, each point indicates the best possible logical error rate for each architecture as we increase the number of transmons, where at each point we pick the optimal code distance (and erasure fraction, for the hybrid architecture) to minimize logical error rate. We find that the best strategy is often to have $100\%$ erasure qubits, but there are ranges where the hybrid approach is potentially better. As the size of the system increases, these regions become proportionally smaller. We observe the same trend in the bottom plot of Figure 12, where for a target logical error rate we plot the fewest number of transmons needed for each architecture.
@@ VI-A Empirical effect of erasure qubit placement
Figure 13 shows logical error rate correlations for each qubit location in a $d=7$ surface code. Samples are generated from circuit-level simulations of randomly distributed erasure qubits of a given erasure fraction. As expected, we generally see that qubits near the center have the highest correlations with the logical error rate, since these qubits are included in the most error paths that connect opposite boundaries. Interestingly, we observe “checkerboard” patterns for independent $Z$ and $X$ logical error rates; we hypothesize that these are an artifact of the different CNOT orderings for $Z$/$X$ stabilizers. Combining the $Z$ and $X$ error rates into a single logical error rate recovers the expected behavior from the code-capacity analysis, where distance from the center is the most important feature.
@@ VI-B Other erasure configurations
First, we show that orientation of the lines do not matter as long as they achieve the same $k_{e}$. We show the $d_{\text{eff}}$ results obtained via simulation in 15 from three such configurations, where lines of erasures take different orientations, where the impact on effective distance is small.
Then, we see that it’s better to place lines of earsures closest to the center. As shown in Figure 16, we see that for an equal number of lines of erasures ($\frac{d+1}{2}$), placing them consecutively in the center yields a higher effective distance than placing them alternately. This follows from our results in Figures 6 and 13, which show that we should expect the middle data qubits to have a higher impact on performance than those on the edge, as prescribed by the optimized placement strategy.

### v2

@@ Abstract
Dual-rail erasure qubits can substantially improve the efficiency of quantum error correction, allowing lower error rates to be achieved with fewer qubits, but each erasure qubit requires $3\times$ more transmons to implement compared to standard qubits. In this work, we introduce a hybrid-erasure architecture for surface code error correction where a carefully chosen subset of qubits is designated as erasure qubits while the rest remain standard. Through code-capacity analysis and circuit-level simulations, we show that a hybrid-erasure architecture can boost the performance of the surface code—much like how a game of Minesweeper becomes easier once a few squares are revealed—while using fewer resources than a full-erasure architecture. We study strategies for the allocation and placement of erasure qubits through analysis and simulations. We then use the hybrid-erasure architecture to explore the trade-offs between per-qubit cost and key logical performance metrics such as threshold and effective distance in surface code error correction. Our results show that the strategic introduction of dual-rail erasure qubits in a transmon architecture can enhance the logical performance of surface codes for a fixed transmon budget, particularly for near-term-relevant transmon counts and logical error rates.
@@ I Introduction
Figure 1: Viewing a hybrid-erasure architecture as a game of Minesweeper. Top: In a standard surface code, we receive syndrome bits (red and blue 1s) indicating that an error string terminates nearby. The decoder must find a valid explanation (red and blue flags) based only on this information. If the decoder guesses wrong, a logical error can occur. Bottom: When data qubits are erasure qubits, we gain extra information before decoding (black bombs and green checks on left) that help us decide which error string to predict, avoiding logical errors and achieving better error suppression. Middle: Partial improvements over error suppression persists when a carefully selected a subset of qubits are erasure qubits.
@@
While typical noise models used to simulate and evaluate QEC assume that physical one- and two-qubit errors are random, uniformly distributed Pauli operators, significant gains in QEC efficiency can be found under alternative noise models [8, 9, 10, 11, 12, 13]. For example, under a biased noise model where the relative strengths of Pauli X and Z errors are orders of magnitude different, the XZZX surface code can drastically improve the noise threshold compared to the standard surface code [8]. Another type of non-standard error model is heralded erasure, in which the information in a physical qubit is completely lost, but the time and location of the error are known. A surface code made up of erasure qubits, in which the primary error mechanism is heralded erasure, can have a higher threshold and, crucially, twice the effective distance: a surface code patch with $d\times d$ data qubits can correct d1d-1 heralded erasure errors on ideal erasure qubits, but only $\lfloor(d-1)/2\rfloor$ errors under a standard Pauli error model. This gives erasure qubits a significant scaling advantage over standard unbiased qubits, leading to significant interest in the proposal [14, 15, 16, 17] and realization [18, 19, 20, 21, 22, 23, 24, 25, 26] of erasure qubits in various hardware modalities.
Dual-rail erasure qubits in superconducting transmons [16, 15] are a particularly promising approach to implementing erasure qubits, with several recent experimental realizations [22, 20, 26]. A dual-rail qubit is encoded in two transmons such that a relaxation error (the dominant error in individual transmons) in either transmon takes the system out of the qubit space; an erasure check can be performed regularly to determine whether this has happened, and if so, the qubit is reset and the information is sent to the decoder. Some dual-rail implementations require a third transmon to perform this erasure check, bringing the relative cost of a single dual-rail erasure qubit to $3\times$ that of a standard qubit.
@@ II-B Erasure qubits
$|\overline{0}\rangle=\frac{|01\rangle+|10\rangle}{\sqrt{2}}\hskip 6.0pt,\hskip 18.0pt|\overline{1}\rangle=\frac{|01\rangle-|10\rangle}{\sqrt{2}}$ (1)
@@ III Analysis of a hybrid-erasure surface code
  • $P\subset[d^{2}]$ specifies the subset of data qubits (indexed by $d^{2}]=\{1,..,d^{2}\}$) designated as erasures.
@@ III-A Summary of results
Figure 4: Examples of hybrid-erasure architecture. (a) $\mathcal{A}(5,0.6,P_{r})$ where the placement $P_{r}=\{1,2,5,...,22\}$ is sampled at random. (b) $\mathcal{A}(7,0.57,P^{*})$. Since $f_{e}\times d^{2}=27.93$, only 2727 erasures are allocated, with $24$ erasures placed in 22 rows/columns, and the remaining 33 placed greedily, all as close to the center as possible.
@@ III-B The repetition code: a toy model
Figure 5: Error correction on the repetition code with heralded erasure noise. (a) We consider a distance-$7$ repetition code with $k=3$ erasure qubits, at indices $q_{2},q_{3},q_{6}$. (b) Suppose we encounter 33 physical $X$ errors, triggering a detection event on $q_{3}$. (c) A physical error on $q_{2}$ may trigger the detection event, but does not flip any syndromes; this could be due to the deplorizing error being projected onto the code space as a $Z$ error or an identity, each with probability $25\%$. (d) There are two ways to match the syndromes. Since there’s no detection even on $q_{6}$, we know the error must be the error chain not containing $q_{6}$.
@@
$\times p^{l_{d}}(1-p)^{d-k-l_{d}}(\frac{1}{2})^{k}$
$\times\chi[2l_{d}<d-k]$
@@ III-C Impact of erasure placement on logical error rate
We show in Figure 6 a histogram of how frequently each qubit in a distance $7$ surface code is contained in a logical error. Immediately, we observe that the qubits in the middle are included in up to $3\times$ more paths than on the edge. We derive our closest-to-center placement strategy from this observation. Secondly, a more subtle point is that, since each lattice path either must traverse all columns or all rows, filling an additional row and column with erasures necessarily guarantees that each lattice path encounters an additional erasure qubit. The effective distance of the code will be guaranteed to increase by the same analysis given in Section III-B. It is important to note that, while in general we should place erasures where they encounter more traversing paths (in the center), we should not prioritize this over filling additional columns/rows, since, to leading order, the logical error performance is constricted by the error with the shortest fault distance.
@@ IV-A Noise model
Initialization error pp $2p$
Readout error pp $2p$
@@
Figure 8: Error suppression characteristics for hybrid erasure architectures via circuit level simulation. Left: Effective distance $d_{\text{eff}}$ of surface codes at various erasure qubit fractions. Optimized placement yields a higher effective distance for intermediate erasure fractions. Lines show linear interpolation between $(d+1)/2$ and $d$. Right: Extracted surface code thresholds for random and optimized placement for different erasure fractions. Optimized placement leads to more consistent improvements in threshold.
@@ IV-B Simulation and decoding
We perform circuit-level simulations using Stim[47]. Errors are added upon initialization, readout, single-qubit gates, two-qubit gates, and erasure checks according to the noise model in Table I. One- and two-qubit gate errors on erasure qubits are modeled as heralded detections followed by a replacement of the qubit(s) with a maximally mixed state. The resulting circuit-level effect is an erasure flag along with $\{I,X,Y,Z\}$ errors with probability $\frac{1}{4}$ for each affected qubit. In Stim, the erasure flag is
@@
where $p_{e}$ is the probability of error $e$ derived from the noise model in Table I.
@@ V Evaluating hybrid-erasure architectures
Figure 9: Achievable logical error rate for hybrid-erasure surface codes as a function of transmon count for several fixed physical error rates, where expected yield (background gradient) is calculated assuming a defect rate of $10^{-2}$ per transmon. Hybrid architectures outperform all-standard or all-erasure architectures for certain transmon budgets.
@@ V-A Effective distance and threshold
Figure 10: Scaling of costs in Figure 9 for larger systems using an all erasure, no erasure, or hybrid erasure architectures. All hybrid points sweep over and select the best erasure fraction. Shaded region represents area of possible hybrid erasure values in between analytical upper bound and approximate observed circuit-level lower bound. Left: Minimum possible logical error rate for fixed transmon cost. Background gradient indicates resulting single chip yield for a defect rate of $10^{-3}$. Right: Minimum possible number of transmons for a target logical error rate.
@@
Figure LABEL:fig:thresholds shows the extracted surface code thresholds for varying erasure qubit fraction, similarly showing a steady improvement as we add more erasure qubits. Again, a random placement of erasure qubits significantly underperforms compared to the optimized heuristic.
@@ V-B Comparing transmon counts
In this context, we examine the transmon cost of hybrid-erasure architectures in Figure 9 for several fixed physical error rates. These can be viewed as vertical slices through Figure 7, where the cost $C(\mathcal{A}(d,f_{e}))=d^{2}\times[(1-f_{e})+3\times f_{e}]$ calculates the number of transmons needed to implement a surface code of distance $d$ and erasure fraction $f$. For a fixed code distance and physical error rate, we see the expected increase in transmon count and improvement in logical error rate as erasure fraction increases. For certain fixed transmon budgets, e.g. from 70 to 100 or from 150 to 190 transmons, a hybrid-erasure design gives a better logical error rate than can be achieved with a full-erasure design.
@@ V-C Scaling to larger transmon counts
Although Figure 9 helps analyze costs for near-term system sizes, we are also interested in the impact of hybrid erasure architectures for larger systems. To estimate the performance of the hybrid erasure architecture in higher distances and lower logical error rates, we use the empirical thresholds and a projected $d_{\text{eff}}$ bounded between two estimates. As an optimistic upper bound, we may expect the erasure fraction to behave as in the repetition code described in Section III-B. We observe that the fit becomes more accurate as we increase the distance of the surface code. The analytical lower-bound is derived using Equation 2. In Figure 10, we report the relative costs and performances of full-erasure, no-erasure, and hybrid architectures implementable over up to 1,000 transmons and down to logical error rates of $10^{-15}$. For each physical error rate, we estimate the logical error rates and the costs achieved by all possible erasure architectures, plotting the best logical error rates achieved by architectures implementable over fixed transmon budgets (Top, Figure 9) and the fewest transmons used achieving fixed target logical error rates (Bottom, Figure 9). The shaded regions represent the parameter regimes where a hybrid architecture may potentially outperform the standard approaches. We find that the best strategy is often to have $100\%$ erasure qubits, but there are ranges where the hybrid approach is potentially better. As the size of the system increases, these regions become proportionally smaller. We observe the same trend in the bottom plot of Figure 10.
@@ VI-A Empirical effect of erasure qubit placement
Figure 11 shows logical error rate correlations for each qubit location in a $d=7$ surface code. Samples are generated from circuit-level simulations of randomly distributed erasure qubits of a given erasure fraction. As expected, we generally see that qubits near the center have the highest correlations with the logical error rate, since these qubits are included in the most error paths that connect opposite boundaries. Interestingly, we observe “checkerboard” patterns for independent $Z$ and $X$ logical error rates; we hypothesize that these are an artifact of the different CNOT orderings for ZZ/$X$ stabilizers. Combining the $Z$ and $X$ error rates into a single logical error rate recovers the expected behavior from the code-capacity analysis, where distance from the center is the most important feature.
@@ VI-B Other erasure configurations
First, we show that orientation of the lines do not matter as long as they achieve the same $k_{e}$. We show the $d_{\text{eff}}$ results obtained via simulation in 13 from three such configurations, where lines of erasures take different orientations, where the impact on effective distance is small.
Then, we see that it’s better to place lines of earsures closest to the center. As shown in Figure 14, we see that for an equal number of lines of erasures ($\frac{d+1}{2}$), placing them consecutively in the center yields a higher effective distance than placing them alternately. This follows from our results in Figures 6 and 11, which show that we should expect the middle data qubits to have a higher impact on performance than those on the edge, as prescribed by the optimized placement strategy.
@@ Acknowledgements
This work is funded in part by the STAQ project under award NSF Phy-232580; in part by the US Department of Energy Office of Advanced Scientific Computing Research, Accelerated Research for Quantum Computing Program; and in part by the NSF Quantum Leap Challenge Institute for Hybrid Quantum Architectures and Networks (NSF Award 2016136), in part based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, and in part by the Army Research Office under Grant Number W911NF-23-1-0077. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work was completed in part with resources provided by the Research Computing Center of the University of Chicago.
FTC is the Chief Scientist for Quantum Software at Infleqtion and an advisor to Quantum Circuits, Inc.